OpenStudy (anonymous):

Tracy ran for seven-twelfths hours and Reina ran for one-third hours. How many more hours did Tracy run? one-fourth 1 one-third one-sixth

OpenStudy (anonymous):

@mkerr017

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

one fourth, just make the two fractions have the same denominator, then subtract the one from the other and simplify it :)

OpenStudy (anonymous):

answer plz

OpenStudy (anonymous):

I just told you, one fourth

OpenStudy (anonymous):

thx \

OpenStudy (texaschic101):

7/12 - 1/3 = (common denominator is 12) 7/12 - 4/12 = 3/12 = 1/4 yep..thats it :)